OpenStudy (anonymous):

At an interest rate of 4% compounded annually, how long will it take to double the following investments? (Round your answer to the nearest year.) (a) $100 _____ years

OpenStudy (anonymous):

why is it not 25?

OpenStudy (anonymous):

200 = 100(1 + 0.4)^n solve for n

OpenStudy (anonymous):

wait isnt n 1

OpenStudy (anonymous):

(1.04)^n = 2 n ln 1.04 = ln 2 n = ln 2/ ln 1.04 = 17.67 18 years

OpenStudy (anonymous):

why does 1.04^n = 2?

OpenStudy (anonymous):

(1.04)^n = 200/100

OpenStudy (anonymous):

i dont get what u did after that. could u retricewords to explain ur steps?

OpenStudy (anonymous):

retrice? i said USE

OpenStudy (anonymous):

have you seen the formula before ie amount = principap ( 1+ annual rate) ^ number of years A = P(1 + r)^n

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

u get 2=1.04^n and then what

OpenStudy (anonymous):

how do u get 18?

OpenStudy (anonymous):

(1.04)^n = 2 when the unknown is the exponent you then take logarithms (logs) of both sides do you know about logs?

OpenStudy (anonymous):

sorry i have to go out for a while marion - maybe we can talk later

OpenStudy (anonymous):

not really no. and thta's fine. the name is mario by the way.

OpenStudy (anonymous):

*that's